OpenStudy (anonymous):

What is the domain and range of f(x)=2x-1 using inequalities and interval notation? x values -4,-3,-2,-1,0,1,2,3,4 y values -9,-7,-5,-3,-1,1,3,5,7

OpenStudy (anonymous):

Domain is all the values of x in the closed interval i:e D is [-4;4] because the function is defined for all the values within that closed interval. Range is all the values between and including -9 and 7 i:e R [-9;7]

OpenStudy (anonymous):

I need to type a paper on this and I am not sure about if it means using arrows with the different brackets was what it means when it says this? using inequalities and interval notation?

OpenStudy (anonymous):

ohw if you use that notation, then your domain will be \[-4 \le x \le 4\] and your range will be \[-9 \le y \le 7\]. Was that helpful?

OpenStudy (anonymous):

Yes thank you so much I wasnt sure if I needed to draw something like this  if so what would you suggest on the brackets?

OpenStudy (anonymous):

anyone with the line and the correct brackets? I want to cover all bases LOL